OpenStudy (anonymous):

help!!! the formula for the perimeter P of a rectangle with length 1 and width w is P=21+2w. what is the formula for the length of a rectangle in terms of the perimeter and width?

OpenStudy (anonymous):

@KeganMoore

OpenStudy (anonymous):

I'm not that good with math. History and English are my zones. Sorry :/

OpenStudy (anonymous):

its ok thnx though.... do u know anyone who can help? @KeganMoore

OpenStudy (anonymous):

If she will respond to a message!!! @Hope_Nicole

OpenStudy (anonymous):

P=2l+2w 2l=p-2w l=\[\frac{ p-2w }{ 2 }\]

OpenStudy (anonymous):

thnx!! everyone